# OmniJARVIS: Unified Vision-Language-Action Tokenization Enables Open-World Instruction Following Agents

### v1

@@ Appendix A Training Details
We utilized the SFTTrainer class from the TRL library by Hugging Face to train the VLM model. The learning rate was set at 1.4e-5, and a cosine learning rate scheduler was employed. The weight decay parameter was set to 0 with a warm-up ratio of 0.03. Training took place on 8 A800 GPUs with FSDP, with a batch size of 2 and gradient accumulation steps of 4 using bf16 precision. The training lasted for one epoch on our generated dataset. The raw interaction dataset comes from the sections 6xx, 7xx, and 10xx of the contractor dataset provided by OpenAI (Baker et al., 2022) and the recording interactions of JARVIS-1 Agents (Wang et al., 2023b).
@@
Each frame in our experiments has a resolution of 128x128 pixels. We segmented each episode into multiple trunks, with each trunk consisting of 128 frames. The learning rate was set at 0.00004, with a weight decay of 0.001. The batch size was configured to 2, and training was conducted on a cluster of eight NVIDIA 3090 Ti graphics cards. The training dataset comprised sections 6xx, 7xx, 9xx, and 10xx of the contractor dataset provided by OpenAI (Baker et al., 2022). The precision for training was set to bfloat16.
@@ Appendix B FSD Computation
This section outlines the computation of FSD score between the generated videos and human gameplay recordings. First, we divide the videos into trunks of 128 frames. For each segment, we sample 16 frames, with 8 frames in between each sampled frame. These sequences of 16 frames are then fed through the video encoder of MineCLIP (Fan et al., 2022) to obtain 512-dimensional video embeddings. Finally, the score is calculated according to (Heusel et al., 2017) between the embeddings of the generated videos and the reference videos.
@@
Appendix C Embodied Question Answering Benchmarks
@@
To enhance the LLM-based agent’s understanding of Minecraft, we gathered a large dataset called Minecraft-Embodied-Question-Answering-300k (mc-eqa-300k). Initially, we extracted language corpus from the Minecraft Wiki [^1] https://minecraft.wiki/ and DigMinecraft [^2] https://www.digminecraft.com/ website to create prompts for the language model (ChatGPT). This resulted in around 50k question-answer pairs. Using these as seed questions, ChatGPT generated approximately 250k diverse questions on various aspects of Minecraft. To ensure accurate responses and avoid hallucinations common with existing models like self-instruct (Wang et al., 2022), we employed a retrieval augmented generation (Lewis et al., 2020) method known as RAT (Wang et al., 2024) to generate accurate answers. The resulting question-answering dataset, along with interaction data, form our training dataset for OmniJARVIS.
@@
Appendix D Prompts for Creative Tasks
Prompts for Creative Tasks 1. Cook the beef with a furnace and recycle the furnace. If you meet night, place and use the bed for sleeping. 2. Explore caves, mine resources, and craft items in Minecraft to progress and survive. 3. Gather resources, craft tools, and cook food in Minecraft. 4. Place a torch on the wall to light the environment. Collect and picking it up when you leave. 5. Craft an oak boat and use it to travel 6. Obtain resources for building and survival by gathering materials and farming resources. 7. Consistently interact with chests to manage inventory contents. 8. Explore and gather resources in Minecraft. 9. Collect and mine azure bluets, deal damage to mobs, and defeat mobs in the game. 10. Do the following tasks sequentially: 1. Gather oak logs and oak leaves from trees. 2. Use oak logs to create oak planks and then a crafting table. 3. Create sticks from oak planks using the crafting table. 4. Craft a wooden axe and a wooden pickaxe using sticks and oak planks. 5. Collect materials like mushrooms and brown mushrooms by mining blocks with the wooden axe. 6. Mine grass, tall grass, and stone using the wooden tools for resources. 11. Harvest sugar cane to obtain multiple sugar cane pieces. 12. Plant and consume wither roses repeatedly. 13. Harvest wheat seeds, plant them, and use the harvested wheat seeds to feed animals or craft items such as bread. 14. Trade with a villager by giving emeralds and books to receive enchanted books as well as new emeralds and books. 15. Mine ice using an iron pickaxe and pick up the ice block obtained. 16. Open a chest in the game to access or manage inventory items.
Appendix E Prompt for Instruction Generation
Prompt E.1:Prompt for Instruction Generation **Instruction**:
@@ Appendix E Prompt for Instruction Generation
Appendix F Prompt for Thought Generation
Prompt F.1:Prompt for Thought Generation A player is playing Minecraft. I need you to give thought about what behavior it should take next given current situation. Here are some demonstrations:
@@ Appendix F Prompt for Thought Generation
Appendix G Prompt for Memory Summarization
Prompt G.1:Prompt for Memory Summarization A player is playing Minecraft.The situation of the player contains 4 parts: task, state, inventory and memory. Under this situation, the player will take a behavior. And after this behavior, the player’s memory will be updated to "Updated Memory". I need you to give a subpart of the player’s updated memory that is most relevant to its task.
@@ Appendix G Prompt for Memory Summarization
Appendix H Examples of OmniJARVIS Interaction Process

### v2

@@ Appendix A Training Details
We utilized the SFTTrainer class from the TRL library by Hugging Face to train the VLM model. The learning rate was set at 1.4e-5, and a cosine learning rate scheduler was employed. The weight decay parameter was set to 0 with a warm-up ratio of 0.03. Training took place on 8 A800 GPUs with FSDP, with a batch size of 2 and gradient accumulation steps of 4 using bf16 precision. The training lasted for one epoch on our generated dataset. The raw interaction dataset comes from the sections 6xx, 7xx, and 10xx of the contractor dataset provided by OpenAI [2] and the recording interactions of JARVIS-1 Agents [47].
@@
Each frame in our experiments has a resolution of 128x128 pixels. We segmented each episode into multiple trunks, with each trunk consisting of 128 frames. The learning rate was set at 0.00004, with a weight decay of 0.001. The batch size was configured to 2, and training was conducted on a cluster of eight NVIDIA 3090 Ti graphics cards. The training dataset comprised sections 6xx, 7xx, 9xx, and 10xx of the contractor dataset provided by OpenAI [2]. The precision for training was set to bfloat16.
@@ Appendix B FSD Computation
This section outlines the computation of FSD score between the generated videos and human gameplay recordings. First, we divide the videos into trunks of 128 frames. For each segment, we sample 16 frames, with 8 frames in between each sampled frame. These sequences of 16 frames are then fed through the video encoder of MineCLIP [16] to obtain 512-dimensional video embeddings. Finally, the score is calculated according to [19] between the embeddings of the generated videos and the reference videos.
@@
Appendix C Benchmarks Details
C.1 Programmatic Tasks
Table 6: Description and setting of programmatic tasks.
Groop
Task
Number
Maximum
Steps
Task
Prompt
Item List
Wooden 10 3000 Make {item} from empty inventory. stick, crafting_table, chest, ladder, bowl, button, door, boat, …
Food 5 6000 Get the food {item}. cooked_chicken, cooked_mutton, cooked_porkchop, cooked_beef, bread
Stone 5 3000 Craft a {item}. charcoal, smoker, stone_sword, furnace, torch
Iron 5 6000 Smelt iron ingots and craft {item}. iron_sword, iron_ingot, bucket, iron_nugget, shears
Diamond 5 12000 Dig down and craft {item}. diamond_pickaxe, diamond_shovel, diamond_hoe, diamond_axe, diamond
C.2 Embodied Question Answering Benchmarks
@@
To enhance the LLM-based agent’s understanding of Minecraft, we gathered a large dataset called Minecraft-Embodied-Question-Answering-300k (mc-eqa-300k). Initially, we extracted language corpus from the Minecraft Wiki [^1] https://minecraft.wiki/ and DigMinecraft [^2] https://www.digminecraft.com/ website to create prompts for the language model (ChatGPT). This resulted in around 50k question-answer pairs. Using these as seed questions, ChatGPT generated approximately 250k diverse questions on various aspects of Minecraft. To ensure accurate responses and avoid hallucinations common with existing models like self-instruct [45], we employed a retrieval augmented generation [26] method known as RAT [48] to generate accurate answers. The resulting question-answering dataset, along with interaction data, form our training dataset for OmniJARVIS.
@@
C.3 Open-ended Instruction-Following Tasks
Prompts for Open-ended Instruction-following Tasks 1. Cook the beef with a furnace and recycle the furnace. If you meet night, place and use the bed for sleeping. 2. Explore caves, mine resources, and craft items in Minecraft to progress and survive. 3. Gather resources, craft tools, and cook food in Minecraft. 4. Place a torch on the wall to light the environment. Collect and picking it up when you leave. 5. Craft an oak boat and use it to travel 6. Obtain resources for building and survival by gathering materials and farming resources. 7. Consistently interact with chests to manage inventory contents. 8. Explore and gather resources in Minecraft. 9. Collect and mine azure bluets, deal damage to mobs, and defeat mobs in the game. 10. Do the following tasks sequentially: 1. Gather oak logs and oak leaves from trees. 2. Use oak logs to create oak planks and then a crafting table. 3. Create sticks from oak planks using the crafting table. 4. Craft a wooden axe and a wooden pickaxe using sticks and oak planks. 5. Collect materials like mushrooms and brown mushrooms by mining blocks with the wooden axe. 6. Mine grass, tall grass, and stone using the wooden tools for resources. 11. Harvest sugar cane to obtain multiple sugar cane pieces. 12. Plant and consume wither roses repeatedly. 13. Harvest wheat seeds, plant them, and use the harvested wheat seeds to feed animals or craft items such as bread. 14. Trade with a villager by giving emeralds and books to receive enchanted books as well as new emeralds and books. 15. Mine ice using an iron pickaxe and pick up the ice block obtained. 16. Open a chest in the game to access or manage inventory items.
Appendix D Prompt for Instruction Generation
Prompt D.1:Prompt for Instruction Generation **Instruction**:
@@ Appendix D Prompt for Instruction Generation
Appendix E Prompt for Thought Generation
Prompt E.1:Prompt for Thought Generation A player is playing Minecraft. I need you to give thought about what behavior it should take next given current situation. Here are some demonstrations:
@@ Appendix E Prompt for Thought Generation
Appendix F Prompt for Memory Summarization
Prompt F.1:Prompt for Memory Summarization A player is playing Minecraft.The situation of the player contains 4 parts: task, state, inventory and memory. Under this situation, the player will take a behavior. And after this behavior, the player’s memory will be updated to "Updated Memory". I need you to give a subpart of the player’s updated memory that is most relevant to its task.
@@ Appendix F Prompt for Memory Summarization
Appendix G Examples of OmniJARVIS Interaction Process